# Interstellar Medium Modulation of Nonlinear Kinetic Alfvén Morphology in Structured Galactic Environments

### v1

@@ Abstract
We present a spatially dependent framework for the existence and propagation of nonlinear kinetic Alfvén (KA) structures in the interstellar medium (ISM). Using a multi-component analytical model that incorporates the diffuse warm ionized medium together with localized H II regions, supernova remnants (SNR), and stellar-wind bubbles (SWB), we derive location-dependent coefficients governing KA dispersion and nonlinearity. The reductive perturbation method is applied to obtain Korteweg-de Vries (KdV) equations, enabling the characterization of solitons under realistic astrophysical conditions. Numerical analysis demonstrates how superthermality, plasma $\beta$, temperature, and density gradients modulate soliton amplitude, width, and stability. Our results reveal distinct exclusion zones (EZs) for KA solitons in high-$\beta$ H II regions and SWB/SNR interiors, as well as ultra low-$\beta$ regions near central pulsar wind nebulae. While H II regions exhibit simple Gaussian-driven depletions, the complex “hole-and-shell” morphologies of SWBs and SNRs imprint sharp spatial variations and discontinuities on soliton properties. This study establishes a direct link between macroscopic ISM morphology, ion-kinetic scale dissipation, and the emergence of coherent Alfvénic activity, with implications for radio scattering, pulsar scintillation, and fine-scale signatures in astrophysical observations.
Interstellar medium—Kinetic Alfvén Waves—Solitons—H II Regions—Stellar-Wind Bubbles—Supernova Remnants.
show]singhmanpreet185@gmail.com
show]liusm@swjtu.edu.cn
I Introduction
@@
At sub-ion scales, kinetic theory predicts that the Alfvénic cascade transitions naturally into kinetic Alfvén waves (KAWs), which mediate energy transfer to small scales and ultimately to particle heating and acceleration. Gyrokinetic theory (Schekochihin et al., 2009) and solar-wind simulations (Howes et al., 2008b, a; Zhao et al., 2016) demonstrate that KAWs dominate between ion and electron gyroscales, with dissipation primarily via Landau damping, and recent work shows that KAW intermittency can lead to efficient electron heating (Zhou et al., 2023). Observations indicate that plasma beta in the WIM spans a broad range (Ferrière, 2020). While some regions approach $\beta\sim 1$, substantial regions of the diffuse ionized medium exhibit $\beta<1$, falling within the regime $m_{e}/m_{i}\ll\beta\ll 1$ (with $m_{e}$ and $m_{i}$ being the electron and ion mass respectively) where KAWs and their nonlinear extensions can exist. Since the WIM is a magnetized, weakly collisional plasma with a cascade extending to ion-kinetic scales, KAW turbulence is a natural candidate for shaping small-scale structure and dissipation in the ISM.
@@
The ISM is highly non-uniform, comprising structures such as H II regions, supernova remnants (SNRs), stellar-wind bubbles (SWBs) etc., each producing sharp gradients in density, temperature, magnetic field strength, and plasma beta (Ferrière, 2001; Elmegreen and Scalo, 2004; Mac Low and Klessen, 2004). Stellar feedback injects energy on parsec scales through supernova explosions, expanding H II regions, and SWBs, maintaining a mildly supersonic, magnetized turbulent cascade across the WIM (Hill et al., 2012). Furthermore, the interiors of SNRs, particularly near Pulsar Wind Nebulae (PWNe), are enriched with relativistic pairs, creating a multi-species electron-positron-ion (although with race ions) plasma that fundamentally alters wave dispersion. Modern numerical studies indicate that the ISM often operates in a strong-turbulence regime where magnetic fluctuations reach $|\delta\mathbf{B}|\gtrsim|\mathbf{B}_{0}|$ (Goldreich and Sridhar, 1995; Perez and Boldyrev, 2008; Schekochihin, 2022), and galaxy-scale simulations show that such large-amplitude disturbances are ubiquitous and driven by stellar feedback (Ntormousi et al., 2024). Strong turbulence naturally generates intermittent coherent structures, such as current sheets, filaments, vortices, and shocklets, that localize dissipation and promote nonlinear wavefront steepening (Uritsky et al., 2010). Such structures are well known to excite finite-amplitude Alfvénic and kinetic-Alfvénic fluctuations (Karimabadi et al., 2013), providing physically motivated seed perturbations that can steepen into localized kinetic Alfvén (KA) solitons under appropriate plasma conditions (Singh et al., 2019). While current sheets, shocklets, and magnetic filaments are well-established products of strong turbulence, the possibility that the same turbulent dynamics can generate KA solitons has not yet been explored in the context of the ISM. If present, KA solitons can produce localized density and magnetic-field fluctuations that influence observables such as radio-wave scattering, pulsar scintillation, or microstructures in timing data, providing potential astrophysical signatures of coherent KA activity.
@@
Because the ISM contains strong spatial variations in magnetic field, density, temperature, and particle composition, the coefficients governing KA dispersion and nonlinearity can vary significantly from region to region. Even modest environmental changes can modify soliton amplitude, width, or existence conditions (Hasegawa and Chen, 1976). Thus, the structured nature of the ISM, including high-$\beta$ H II cavities, compressed SNR shells, and density-evacuated SWBs can directly influence the formation, propagation, and characteristics of nonlinear KAW solitons (Subramanian et al., 2006; Kivotides et al., 2007; Kritsuk et al., 2017; Richard et al., 2022; Colman et al., 2022; Galishnikova et al., 2022; Lübke et al., 2024; Lesaffre et al., 2024). In addition to their spatial variability, fine-scale KA structures are efficient sites of electron heating in weakly collisional plasmas (TenBarge et al., 2013), indicating that the soliton structures considered here may influence local ISM thermodynamics.
@@
The velocity distribution of electrons in the ISM can deviate from a strict Maxwellian, particularly in weakly collisional plasmas subject to continuous driving and intermittent energization. These conditions produce suprathermal power-law tails, which are often modeled using the $\kappa$-distribution, a framework originally developed for space plasma contexts and now standard for characterizing non-thermal populations (e.g., Vasyliunas, 1968; Summers and Thorne, 1991; Pierrard and Lazar, 2010; Livadiotis and McComas, 2009). In this formulation, the $\kappa$-index quantifies the departure from thermal equilibrium: a smaller $\kappa$ corresponds to a harder high-energy tail, while the Maxwellian distribution is smoothly recovered as $\kappa\rightarrow\infty$ (in practice, $\kappa\gtrsim 100$ is often indistinguishable from a Maxwellian). Within ISM context, $\kappa$-distributed electrons have been proposed for photoionized nebulae (H II regions and planetary nebulae), although this specific application has been challenged by work arguing that the electron energy distribution is extremely close to Maxwellian under typical nebular conditions (Nicholls et al., 2012; Draine and Kreisch, 2018). More broadly in the structured ISM, persistent suprathermal electrons are physically expected where continuous energization competes with thermalization. This includes supernova-driven collisionless shocks and superbubbles, which generate non-thermal tails via diffusive shock acceleration and collective acceleration mechanisms, respectively, as well as turbulence enabled magnetic reconnection, which can efficiently accelerate particles and produce nonthermal spectra (Drury, 1983; Parizot et al., 2004; Lazarian and Vishniac, 1999; Guo et al., 2014). In our KA soliton context, adopting $\kappa$ is therefore not merely phenomenological: it changes the electron kinetic response that closes the parallel electric field and charge-separation physics underlying dispersive KAW dynamics, and thereby provides a physically motivated control on KA soliton morphology most strongly on the soliton width, especially near sharp gradients and kinetic regime boundaries (Masood et al., 2015)
@@
II Theoretical Model
The ISM is a vast, multi-phase plasma environment governed by the gravitational and magnetic fields of the Milky Way. Its inherent inhomogeneity means that the properties of any propagating wave will be strongly modulated by its location and proximity to different galactic structures, including H II regions, SWBs, and SNRs, among others. To model this, we distinguish between the global coordinate system of the galaxy and the local coordinate system in which the wave dynamics are resolved.
@@ II Theoretical Model
II.1 Multi-Component Representation of the ISM
@@
Below, we describe the background ISM and each embedded structure individually, specifying for each its electron density, magnetic field, temperature, and positron fraction (where applicable). The total values used in the soliton analysis are obtained by summing all contributions. We parameterize the model in terms of electron quantities (e.g., density $n_{e,\text{total}}$ and temperature $T_{e,\text{total}}$) because these are the primary observables constrained by astronomical surveys (e.g., pulsar dispersion measures, thermal emission lines). Complementing this, we incorporate a modeled positron fraction ($p_{\text{total}}$) based on standard PWN theories to accurately represent pair-enriched environments where $n_{e,\text{total}}$ significantly exceeds ion density $n_{i,\text{total}}$. However, since KAWs are fundamentally ion-scale modes governed by ion inertia and finite Larmor radius effects, the corresponding ion density $n_{i,\text{total}}$, which sets the local Alfvén speed, is subsequently derived from these electron and positron profiles via charge neutrality.
II.2 Analytical Formalism for Localized Structures
@@
$\mathcal{S}(d;A,R_{\mathrm{sh}},\sigma_{\mathrm{in}},\sigma_{\mathrm{out}})\\ \equiv A\begin{cases}\exp\left[-\dfrac{(R_{\mathrm{sh}}-d)^{2}}{2\sigma_{\mathrm{in}}^{2}}\right],&d\leq R_{\mathrm{sh}},\\[10.0pt] \exp\left[-\dfrac{(d-R_{\mathrm{sh}})^{2}}{2\sigma_{\mathrm{out}}^{2}}\right],&d>R_{\mathrm{sh}},\end{cases}$ (5)
@@
II.3 Physical Modeling of ISM Components
Now, in this section, we construct a spatially dependent, yet analytically controlled model of the ionized ISM that can be used to evaluate the local plasma parameters entering the KA-soliton framework. The large-scale WIM is modeled as a smooth Galactic background, while discrete astrophysical structures are treated as localized perturbations superposed on this background using the analytical profiles introduced in Sec. II.2.
II.3.1 Galactic Background
Electron Density: The background electron density, $n_{e,\mathrm{bkg}}$, represents the diffuse WIM. It is modeled after the NE2001 framework (Cordes and Lazio, 2002), with typical densities of $\sim 0.01$$0.1\,\text{cm}^{-3}$ (Gaensler et al., 2008) sustained by the leakage of far-ultraviolet radiation from the galactic disk (Haffner et al., 2009). The density profile is defined as:
@@ II.3.1 Galactic Background
Positron Fraction: Unlike electrons and ions, which constitute the primordial thermal plasma, positrons are secondary particles produced by energetic processes (Weidenspointner et al., 2008; Prantzos, 2011). While they eventually annihilate, a diffuse population is maintained in dynamic equilibrium by continuous injection from cosmic-ray spallation (Moskalenko and Strong, 1998) and radioactive decay (Knödlseder et al., 2005). We model the total positron fraction as the ratio of local positron density, $n_{p,\text{total}}$ to the local ion density, $n_{i,\text{total}}$ i.e., $p_{\text{total}}\equiv n_{p,\text{total}}/n_{i,\text{total}}$. This is constructed as a linear superposition of a large-scale Galactic background and localized sources. The large-scale component, $p_{\text{large-scale}}(R)$, consists of a small, uniform background, $p_{\text{bkg}}$, representing the steady-state diffuse population, and a centrally concentrated Galactic source, $p_{\text{GC}}(R)$:
@@
Magnetic Field: The Galactic magnetic field is a crucial agent in ISM dynamics, and its structure is similarly multi-scale. The total magnetic field strength is written as a sum of large-scale background component and self-consistent local perturbations from astrophysical structures. The background field, $B_{0,\text{bkg}}$, is based on the disk-halo paradigm, consistent with galactic dynamo theory and extensive observational data from Faraday rotation and synchrotron emission (Jansson and Farrar, 2012). This component represents the large-scale, ordered field generated by the differential rotation of the galactic disk. So the large-scale galactic magnetic field can be written as:
$B_{0,\mathrm{bkg}}(R,Z)=B_{\odot}e^{-(R-R_{\odot})/L_{B}}e^{-|Z|/h_{B,1}}\\ +\frac{B_{\mathrm{halo}}}{1+(|Z|/h_{B,2})^{2}},$ (9)
@@
Temperature: The ISM exists in a state of thermal multiphase equilibrium, where regional temperatures vary significantly depending on the local balance of heating and cooling mechanisms (Ferrière, 2001). Our model for the total plasma temperature, $T_{e,\mathrm{total}}$, reflects this by superposing intense local heating sources onto a cooler, quasi-isothermal background. The background temperature, $T_{e,\mathrm{bkg}}$, is set to the characteristic value of the WIM:
$T_{e,\mathrm{bkg}}\approx 8000\,\mathrm{K}.$
This temperature is maintained by a delicate equilibrium between large-scale, diffuse heating from the Galaxy’s far-ultraviolet radiation field (Haffner et al., 2009) and cooling via atomic forbidden-line emission from ions such as O+ and N+ (Osterbrock and Ferland, 2006).
II.3.2 H II Regions
Electron Density: Superposed on the quiescent background are distinct regions of significantly altered density, which we model as localized perturbations using specific analytical forms. First, H II regions represent strong positive electron density enhancements, $n_{e,\text{H\,II}}$. These are islands of plasma, often visible as bright emission nebulae, generated by intense photoionizing radiation from central O- or B-type massive stars (Haffner et al., 2009). Since these regions form from the gravitational collapse of dense molecular clouds, their internal density is significantly higher than the ambient ISM (Pagel et al., 1979; Dopita and Evans, 1986; Brinchmann et al., 2008; Shirazi et al., 2014; Li et al., 2025). Using the formalism from Sec. II.2, we model this contribution as a localized Gaussian enhancement:
Magnetic Field: In H II regions, the high thermal pressure of the hot, ionized gas drives an outward expansion. This expansion displaces the frozen-in field lines, creating a magnetic cavity with reduced field strength (Spitzer, 1978; Ferrière, 2001). We model this effect as a localized Gaussian depletion:
Temperature: Inside H II regions, photoionization heating, where excess energy from ionizing photons is converted into kinetic energy of electrons, elevates the temperature above the background WIM (Draine, 2011). Crucially, this temperature rise is limited by a thermostat effect, as temperature increases, cooling via collisionally excited forbidden lines rises exponentially, effectively limiting the equilibrium temperature typically near $10^{4}$ K (Osterbrock and Ferland, 2006). We model this local thermal enhancement as:
II.3.3 Stellar-Wind Bubbles
Electron Density: The powerful radiatively driven winds from massive stars excavate large, low-density cavities in the ISM, surrounded by dense shells of swept-up gas (Castor et al., 1975; Weaver et al., 1977). To model this morphology, we construct the electron density profile as the superposition of a central depletion and a shell enhancement:
Magnetic Field: The magnetic structure of the bubble is governed by flux freezing. As the wind bubble inflates, the frozen-in magnetic field lines are stretched and diluted within the volume, while being strongly compressed into the dense shell at the boundary. We model this as:
Electron Temperature: Standard SWB models describe the interior as a region of hot, shocked stellar wind ($T\sim 10^{6}$ K) that maintains high pressure to support the bubble (Weaver et al., 1977). Because the sound speed is high, the interior evolves into an isobaric region with a nearly uniform temperature distribution. To reproduce this characteristic flat-top thermal profile, we model the interior temperature using the super-Gaussian plateau function defined in Eq. 4. The total electron temperature at SWB includes this plateau and the shock-heated shell:
II.3.4 Supernova Remnants
Electron Density: The expanding forward shock from an SNR sweeps up the surrounding interstellar gas, compressing the plasma into a dense, turbulent shell while simultaneously evacuating the interior (Chevalier, 1977; Vink, 2012). This process results in a complex density morphology consisting of a swept-up shell and a depleted cavity. In composite SNRs, this cavity also hosts a central PWN, injected by the relativistic wind of the compact remnant (Kennel and Coroniti, 1984; Gaensler and Slane, 2006).

### v2

@@ Abstract
Context. We present a spatially dependent framework for the existence and propagation of nonlinear kinetic Alfvén (KA) structures in the interstellar medium (ISM).
Aims. By employing a multi-component analytical model, we aim to derive location-dependent coefficients governing KA dispersion and nonlinearity across structured Galactic environments.
Methods. The model incorporates the diffuse warm ionized medium together with localized H II regions, supernova remnants (SNR), and stellar-wind bubbles (SWB). The reductive perturbation method is applied to obtain Korteweg-de Vries (KdV) equations, enabling the characterization of solitons under realistic astrophysical conditions while accounting for superthermality, plasma beta $(\beta)$, temperature, and density gradients.
Results. Our results reveal distinct exclusion zones (EZs) for KA solitons in high-$\beta$ H II regions and SWB/SNR interiors, as well as ultra low-$\beta$ regions near central pulsar wind nebulae. While H II regions exhibit simple Gaussian-driven depletions, the complex “hole-and-shell” morphologies of SWBs and SNRs imprint sharp spatial variations and discontinuities on soliton properties.
Conclusions. This study establishes a direct link between macroscopic ISM morphology, ion-kinetic scale dissipation, and the emergence of coherent Alfvénic activity. These findings provide a new interpretive framework for radio scattering, pulsar scintillation, and fine-scale signatures in astrophysical observations.
Key Words.:
Interstellar medium – Kinetic Alfvén Waves – Solitons – H II Regions – Stellar-Wind Bubbles – Supernova Remnants.
1 Introduction
@@
At sub-ion scales, kinetic theory predicts that the Alfvénic cascade transitions naturally into kinetic Alfvén waves (KAWs), which mediate energy transfer to small scales and ultimately to particle heating and acceleration. Gyrokinetic theory (Schekochihin et al., 2009) and solar-wind simulations (Howes et al., 2008b, a; Zhao et al., 2016) demonstrate that KAWs dominate between ion and electron gyroscales, with dissipation primarily via Landau damping, and recent work shows that KAW intermittency can lead to efficient electron heating (Zhou et al., 2023). Observations indicate that plasma beta in the WIM spans a broad range (Ferrière, 2001). While some regions approach $\beta\sim 1$, substantial regions of the diffuse ionized medium exhibit $\beta<1$, falling within the regime $m_{e}/m_{i}\ll\beta\ll 1$ (with $m_{e}$ and $m_{i}$ being the electron and ion mass respectively) where KAWs and their nonlinear extensions can exist. Since the WIM is a magnetized, weakly collisional plasma with a cascade extending to ion-kinetic scales, KAW turbulence is a natural candidate for shaping small-scale structure and dissipation in the ISM.
@@
The ISM is highly non-uniform, comprising structures such as H II regions, supernova remnants (SNRs), stellar-wind bubbles (SWBs) etc., each producing sharp gradients in density, temperature, magnetic field strength, and plasma beta (Ferrière, 2001; Elmegreen and Scalo, 2004; Mac Low and Klessen, 2004). Stellar feedback injects energy on parsec scales through supernova explosions, expanding H II regions, and SWBs, maintaining a mildly supersonic, magnetized turbulent cascade across the WIM (Hill et al., 2012). Furthermore, the interiors of SNRs, particularly near Pulsar Wind Nebulae (PWNe), are enriched with relativistic pairs, creating a multi-species electron-positron-ion (although with trace ions) plasma that fundamentally alters wave dispersion. Modern numerical studies indicate that the ISM often operates in a strong-turbulence regime where magnetic fluctuations reach $|\delta\mathbf{B}|\gtrsim|\mathbf{B}_{0}|$ (Goldreich and Sridhar, 1995; Perez and Boldyrev, 2008; Schekochihin, 2022), and galaxy-scale simulations show that such large-amplitude disturbances are ubiquitous and driven by stellar feedback (Ntormousi et al., 2024). Strong turbulence naturally generates intermittent coherent structures, such as current sheets, filaments, vortices, and shocklets, that localize dissipation and promote nonlinear wavefront steepening (Uritsky et al., 2010). Such structures are well known to excite finite-amplitude Alfvénic and kinetic-Alfvénic fluctuations (Karimabadi et al., 2013), providing physically motivated seed perturbations that can steepen into localized kinetic Alfvén (KA) solitons under appropriate plasma conditions (Singh et al., 2019). While current sheets, shocklets, and magnetic filaments are well-established products of strong turbulence, the possibility that the same turbulent dynamics can generate KA solitons has not yet been explored in the context of the ISM. If present, KA solitons can produce localized density and magnetic-field fluctuations that influence observables such as radio-wave scattering, pulsar scintillation, or microstructures in timing data, providing potential astrophysical signatures of coherent KA activity.
@@
Because the ISM contains strong spatial variations in magnetic field, density, temperature, and particle composition, the coefficients governing KA dispersion and nonlinearity can vary significantly from region to region. Even modest environmental changes can modify soliton amplitude, width, or existence conditions. Thus, the structured nature of the ISM, including high-$\beta$ H II cavities, compressed SNR shells, and density-evacuated SWBs can directly influence the formation, propagation, and characteristics of nonlinear KA solitons (Subramanian et al., 2006; Kritsuk et al., 2017). In addition to their spatial variability, fine-scale KA structures are efficient sites of electron heating in weakly collisional plasmas (TenBarge et al., 2013), indicating that the soliton structures considered here may influence local ISM thermodynamics.
@@
The velocity distribution of electrons in the ISM can deviate from a strict Maxwellian, particularly in weakly collisional plasmas subject to continuous driving and intermittent energization. These conditions produce suprathermal power-law tails, which are often modeled using the $\kappa$-distribution, a framework originally developed for space plasma contexts and now standard for characterizing non-thermal populations (e.g., Vasyliunas, 1968; Summers and Thorne, 1991; Pierrard and Lazar, 2010; Livadiotis and McComas, 2009). In this formulation, the $\kappa$-index quantifies the departure from thermal equilibrium: a smaller $\kappa$ corresponds to a harder high-energy tail, while the Maxwellian distribution is smoothly recovered as $\kappa\rightarrow\infty$ (in practice, $\kappa\gtrsim 100$ is often indistinguishable from a Maxwellian). Within ISM context, $\kappa$-distributed electrons have been proposed for photoionized nebulae (H II regions and planetary nebulae), although this specific application has been challenged by work arguing that the electron energy distribution is extremely close to Maxwellian under typical nebular conditions (Nicholls et al., 2012; Draine and Kreisch, 2018). More broadly, in the structured ISM, persistent suprathermal electrons are physically expected where continuous energization competes with thermalization. This includes supernova-driven collisionless shocks and superbubbles, which generate non-thermal tails via diffusive shock acceleration and collective acceleration mechanisms, respectively, as well as turbulence enabled magnetic reconnection, which can efficiently accelerate particles and produce nonthermal spectra (Drury, 1983; Parizot et al., 2004; Lazarian and Vishniac, 1999; Guo et al., 2014). In our KA soliton context, adopting $\kappa$ is therefore not merely phenomenological: it changes the electron kinetic response that closes the parallel electric field and charge-separation physics underlying dispersive KAW dynamics, and thereby provides a physically motivated control on KA soliton morphology most strongly on the soliton width, especially near sharp gradients and kinetic regime boundaries (Masood et al., 2015)
@@
2 Theoretical Model
The ISM is a vast, multi-phase plasma environment governed by the gravitational and magnetic fields of the Milky Way. Its inherent inhomogeneity leads to the modulation of the properties of any propagating wave by its location and proximity to different galactic structures, including H II regions, SWBs, and SNRs, among others. To model this, we distinguish between the global coordinate system of the galaxy and the local coordinate system in which the wave dynamics are resolved.
@@ 2 Theoretical Model
2.1 Multi-Component Representation of the ISM
@@
Below, we describe the background ISM and each embedded structure individually, specifying for each its electron density, magnetic field, temperature, and positron fraction (where applicable). We parameterize the model in terms of electron quantities (e.g., density $n_{e,\text{total}}$ and temperature $T_{e,\text{total}}$) because these are the primary observables constrained by astronomical surveys (e.g., pulsar dispersion measures, thermal emission lines). Complementing this, we incorporate a modeled positron fraction ($p_{\text{total}}$) based on standard PWN theories to accurately represent pair-enriched environments where $n_{e,\text{total}}$ significantly exceeds ion density $n_{i,\text{total}}$. However, since KAWs are fundamentally ion-scale modes governed by ion inertia and finite Larmor radius effects, the corresponding ion density $n_{i,\text{total}}$, which sets the local Alfvén speed, is subsequently derived from these electron and positron profiles via charge neutrality.
2.2 Analytical Formalism for Localized Structures
@@
$\mathcal{S}(d;A,R_{\mathrm{sh}},\sigma_{\mathrm{in}},\sigma_{\mathrm{out}})\equiv A\begin{cases}\exp\left[-\dfrac{(R_{\mathrm{sh}}-d)^{2}}{2\sigma_{\mathrm{in}}^{2}}\right],&d\leq R_{\mathrm{sh}},\\[10.0pt] \exp\left[-\dfrac{(d-R_{\mathrm{sh}})^{2}}{2\sigma_{\mathrm{out}}^{2}}\right],&d>R_{\mathrm{sh}},\end{cases}$ (5)
@@
2.3 Modeling of ISM Components
Now, we construct a spatially dependent model of the ionized ISM that can be used to evaluate the local plasma parameters entering the KA-soliton framework. The large-scale WIM is modeled as a smooth Galactic background, while discrete astrophysical structures are treated as localized perturbations superposed on this background using the analytical profiles introduced in Sec. 2.2.
2.3.1 Galactic Background
The background electron density, $n_{e,\mathrm{bkg}}$, represents the diffuse WIM. It is modeled after the NE2001 framework (Cordes and Lazio, 2002), with typical densities of $\sim 0.01$$0.1\,\text{cm}^{-3}$ (Gaensler et al., 2008) sustained by the leakage of far-ultraviolet radiation from the galactic disk (Haffner et al., 2009). The density profile is defined as:
@@ 2.3.1 Galactic Background
Unlike electrons and ions, which constitute the primordial thermal plasma, positrons are secondary particles produced by energetic processes (Weidenspointner et al., 2008; Prantzos et al., 2011). While they eventually annihilate, a diffuse population is maintained in dynamic equilibrium by continuous injection from cosmic-ray spallation (Moskalenko and Strong, 1998) and radioactive decay (Knödlseder et al., 2005). We model the total positron fraction as the ratio of local positron density, $n_{p,\text{total}}$ to the local ion density, $n_{i,\text{total}}$ i.e., $p_{\text{total}}\equiv n_{p,\text{total}}/n_{i,\text{total}}$. This is constructed as a linear superposition of a large-scale Galactic background and localized sources. The large-scale component, $p_{\text{large-scale}}(R)$, consists of a small, uniform background, $p_{\text{bkg}}$, representing the steady-state diffuse population, and a centrally concentrated Galactic source, $p_{\text{GC}}(R)$:
@@
The Galactic magnetic field is a crucial agent in ISM dynamics, and its structure is similarly multi-scale. The total magnetic field strength is written as a sum of large-scale background component and self-consistent local perturbations from astrophysical structures. The background field, $B_{0,\text{bkg}}$, is based on the disk-halo paradigm, consistent with galactic dynamo theory and extensive observational data from Faraday rotation and synchrotron emission (Jansson and Farrar, 2012). This component represents the large-scale, ordered field generated by the differential rotation of the galactic disk. So the large-scale galactic magnetic field can be written as:
$B_{0,\mathrm{bkg}}(R,Z)=B_{\odot}e^{-(R-R_{\odot})/L_{B}}e^{-|Z|/h_{B,1}}+\frac{B_{\mathrm{halo}}}{1+(|Z|/h_{B,2})^{2}},$ (9)
@@
The ISM exists in a state of thermal multiphase equilibrium, where regional temperatures vary significantly depending on the local balance of heating and cooling mechanisms (Ferrière, 2001). Our model for the total plasma temperature, $T_{e,\mathrm{total}}$, reflects this by superposing intense local heating sources onto a cooler, quasi-isothermal background. The background temperature, $T_{e,\mathrm{bkg}}$, is set to the characteristic value of the WIM: $T_{e,\mathrm{bkg}}\approx 8000\,\mathrm{K}$. This temperature is maintained by a delicate equilibrium between large-scale, diffuse heating from the Galaxy’s far-ultraviolet radiation field (Haffner et al., 2009) and cooling via atomic forbidden-line emission from ions such as O+ and N+ (Osterbrock and Ferland, 2006).
2.3.2 Model of H II Regions
A detailed analytical formulation of the H II region model is provided in Appendix A.
2.3.3 Model of Stellar-Wind Bubbles
The full analytical development of the SWB model is presented in Appendix B.
2.3.4 Model of Supernova Remnants